OpenStudy (alivejeremy):

Simplify. 2m2 – 4n2 + 3m – 6m2 – 2m + 7n2 A. −4m2 + 3n2 + m B. −4m2 + 3n2 C. –36m + 28n D. 9m2 + 6m + 11n2

OpenStudy (alivejeremy):

@Mehek14

OpenStudy (alivejeremy):

i think its B or A

OpenStudy (alivejeremy):

can someone help me

OpenStudy (alivejeremy):

can someone help me

OpenStudy (michele_laino):

hint: I can rewrite the original expression, grouping similar terms, like below: \[\left( {2{m^2} - 6{m^2}} \right) + \left( {7{n^2} - 4{n^2}} \right) + \left( {3m - 2n} \right) = ...?\] please continue

OpenStudy (alivejeremy):

thanks for the hint but i still barly get it

OpenStudy (michele_laino):

oops.. I have made a typo: \[\left( {2{m^2} - 6{m^2}} \right) + \left( {7{n^2} - 4{n^2}} \right) + \left( {3m - 2m} \right) = ...?\]

OpenStudy (michele_laino):

for example, what is: \(2m^2-6m^2=(2-6)m^2=...?\)

OpenStudy (alivejeremy):

i have no idea

OpenStudy (michele_laino):

what is \(2-6=...?\)

OpenStudy (alivejeremy):

4

OpenStudy (michele_laino):

it is \(-4\), since \(-4+6=2\) so the first term is: \(-2m^2\)

OpenStudy (michele_laino):

similarly, what is: \(7n^2-4n^2=(7-4)n^2=...?\)

OpenStudy (michele_laino):

hint: what is \(7-4=...?\)

OpenStudy (alivejeremy):

3

OpenStudy (michele_laino):

perfect! So, the second term is \(3n^2\) finally, what is: \(3m-2m=(3-2)m=...?\)

OpenStudy (michele_laino):

hint: \(3-2=...?\)

OpenStudy (alivejeremy):

1

OpenStudy (michele_laino):

that's right! so the third term is: \(m\), and the final result is: \(-4m^2+3n^2+m\) So, what is the right option?

OpenStudy (alivejeremy):

-1

OpenStudy (michele_laino):

I meant what is the right option: A, B, C or D?

OpenStudy (alivejeremy):

B

OpenStudy (alivejeremy):

?

OpenStudy (michele_laino):

we got the result, from the discussion above, such result is: \[\begin{gathered} \left( {2{m^2} - 6{m^2}} \right) + \left( {7{n^2} - 4{n^2}} \right) + \left( {3m - 2m} \right) = ...? \hfill \\ \hfill \\ = - 4{m^2} + 3{n^2} + m \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, what is the right option: A, B, C or D?

OpenStudy (anonymous):

Medal Michele

OpenStudy (michele_laino):

option B is not the right option, since it contains only two terms, whereas the final result contains three terms

OpenStudy (alivejeremy):

i meant to say A

OpenStudy (michele_laino):

that's right! :)

OpenStudy (anonymous):

Good job! :)

OpenStudy (michele_laino):

thanks!! :) @FireFang

OpenStudy (anonymous):

No problem!